Find a logarithmic equation that relates y and x. (Round all numerical values to three decimal places.)

ln(y) =  

x 1 2 3 4 5 6
y 1 0.630 0.481 0.397 0.342 0.303
Answer #1

Solution :- čo 1 2 3 4 7 6 y 0.630 0.481 0.397 0.342 0.303 en glo -0.4620 -0.731 -0.923 - 1.012 -1.194 - Sun n = [rifn=3.5 ,Trond load Sny = I (mi-n c eny.-eny) = [nilnye-nneng 2 – 19.333 - (-15.3363) =-3.9967 Trend Line: Y=AB² B = emp (saya) = cup
Answer #2
x123456
y1.2501.1461.0901.0511.0220.999